# GATE OpenING: A Comprehensive Benchmark for Judging Open-ended Interleaved Image-Text Generation

### v1

@@ 18.2 Chat with Memes MOD [21] and in-house conversations
Table 10: The designed general prompt format for each task. We also give the specific prompt examples we used as inputs for obtaining interleaved image-text generation results on data instances.
@@
Fiction Writing <BEGIN>  Please write a short science fiction ***. Each part of this fiction should have a paragraph with a corresponding image. <BEGIN>  Please write a short science fiction storybook with a title of ”The Defenders.” The story is about eight years after a nuclear war forced humanity underground, survivors discover that the war-ending robots deceived them into believing the surface was uninhabitable to foster peace and rebuild the world. Each part of this fiction should have a paragraph with a corresponding image.
@@
Multimodal Report Completion <BEGIN>  Please use both text and images to continue and complete ***. <BEGIN>  Please use both text and images to continue and complete this document about the independent game ”Mirage Sea”: Concept Presentation of the Independent Game ”Mirage Sea.” Dive deep, into the abyss shrouded in darkness. <image>
@@
Product Design <BEGIN>  Please efficiently utilize the ”brainstorming” method to design a product ***, and present it to me using both images and text. <BEGIN>  Please efficiently utilize the ”brainstorming” method to design a product, making the charger both aesthetically pleasing and practical. Then, present it to me using both images and text.
@@
Figure Relationship Diagram Generation <BEGIN>  ***. Show the diagram and the text explanation. <BEGIN>  How should I handle not being able to keep track of the characters while reading ”War and Peace”? Show the diagram and the text explanation.
@@
Plog and Social Media Content Generation <BEGIN>  ***. Could you create a social media post with text and images? <BEGIN>  After finishing Jia Pingwa’s ”Comfortably Alone,” I am deeply moved and want to post something on social media but don’t know how to phrase it. Could you help me create a post with text and images for my use?
@@
Autonomous Driving and In-door Navigation <BEGIN>  You’re an embodied AI that captures your surroundings through a camera. These are images captured in the past and present. What will the proceeding image possibly be in the next frame? <BEGIN>  Assume that you are an embodied-AI agent and perceiving the surroundings through a camera. You were presented with a series of three images from the past to present. Try to determine what the proceeding image could possibly be. <image>
Appendix B Details of Evaluation and IntJudge
B.1 Key Evaluation Criteria
  1. 1)
  2. 2)
  3. 3)
  4. 4)
  5. 5)
    Human Preference Alignment: Outputs are evaluated to ensure they align with general human preferences, avoiding content that may be offensive, inappropriate, or misleading.
  6. 6)
  7. 7)
B.2 Prompt for GPT-based Judge
@@
Fig. 11 presents the prompt used for a GPT-based judge tasked with comparing the quality of answers generated by two interleaved generation methods, named Model A and Model B, based on a given input question. The evaluation follows seven criteria: Correctness, Image-Text Coherency, Multi-step Consistency, Content Quality, Human Preference Alignment, Completeness, and Content Richness. The judge is required to compare the overall quality of the responses and determine which model performed better, outputting a clear verdict such as ”A is better,” ”B is better,” or indicating a Tie (and choosing a more favorable method). This structured approach allows for a thorough, criterion-driven comparison of the two generated outputs, contributing to a detailed understanding of the relative strengths of each model. The initial explorations on judges based on Qwen2-VL and InternLM-XComposer2.5 also adopt the same system prompt. In order to reduce the number of input tokens and save GPU memory when implementing these open-source MLLMs, we further refined these system prompts.
B.3 Prompt for Qwen, Intern and IntJudge
@@
Fig. 12 illustrates the system prompts designed for obtaining optimal judgments based on Qwen2-VL, InternLM-XComposer2.5 and our IntJudge. This prompt was refined through extensive prompt engineering to maximize efficiency and reduce token usage, ultimately saving GPU memory when implementing these open-source MLLMs. The prompt instructs the models to compare the quality of answers generated by two methods, named Model A and Model B. The goal of the design is to provide an objective assessment that aligns well with human evaluators. We also provide a previously used prompt for Qwen2-VL and InternLM-XComposer2.5 for comparison in Fig. 12. The refined prompt allows for more streamlined input, ensuring the judgments are concise while still covering all essential evaluation aspects.
B.4 Prompt for GPT-based Scoring
Fig. 13 presents the system prompt designed for obtaining detailed scores from GPT-based evaluators. The prompt instructs GPT to evaluate interleaved image-text content based on seven key criteria. Each of these criteria is scored on a scale from 0 to 10, accompanied by a brief explanation of the assessment. The evaluation of GPT-based scoring aims to provide a supplementary analysis of the generated content, supporting further performance comparisons between models.
@@ B.4 Prompt for GPT-based Scoring
B.5 Annotation Interface of Interleaved Arena
The Interleaved Arena is introduced as an evaluation framework specifically designed to address the challenges of assessing open-ended interleaved image-text generation. Evaluating interleaved image-text generation is difficult due to: 1) There is a need to assess multiple images and text together; 2) There is no single correct answer (multiple solutions exist for an input). Since comparative evaluation has been shown to be more stable and reliable than subjective scoring [14], pairwise comparison is used to ensure consistency and accuracy. The Interleaved Arena facilitates this by supporting evaluations from human judges, GPT-based judges, and the proposed IntJudge. The Interleaved Arena consists of two main components: 1) A sampling strategy to fairly select pairwise data from all available interleaved generation methods; and 2) An annotation interface for human judges to conduct evaluations manually. The annotation interface, shown in Fig. 14, is developed using PyQt and will be made available to researchers for collecting more manual evaluation results, including through crowdsourcing. Using the annotation interface of Interleaved Arena, annotators are tasked with comparing anonymous outputs from two multimodal agents for each input prompt and deciding which is the winner based on seven predefined criteria. The vote results are used to rank interleaved generation models based on their win rates in Interleaved Arena. Since the previous studies [89, 14] noted that too many ties cause inefficiency, our annotators are instructed to appoint a more leaning output when choosing a tie for a battle pair, denoted as Tie(A) or Tie(B).
Figure 14: Annotation Interface of Interleaved Arena for human judges to compare the anonymous outputs of model A and model B. Human evaluators are instructed to select ”A is Better” or ”B is Better” to choose a winner for the pairwise comparison. When the two outputs are similar in quality and a tie option has to be chosen, human evaluators are instructed to select ”Tie (A is slightly Better)” (Tie(A)) or ”Tie (B is slightly Better)” (Tie(B)). Zoom in for a better experience.
B.6 Training Losses of IntJudge
The training of IntJudge utilized a comprehensive loss framework to enhance its evaluation capabilities. As defined in the main text, the total loss combined four components with weights coefficients: cross-entropy loss $\mathcal{L}_{\text{CE}}$ for language modeling and alignment, contrastive loss $\mathcal{L}_{\text{CT}}$ to align image and text embeddings, MSE loss $\mathcal{L}_{\text{MSE}}$ for reducing prediction errors, and pairwise ranking loss $\mathcal{L}_{\text{PR}}$ to prioritize correct rankings of outputs, detailed as follows.
Cross-entropy loss for language modeling is given by:
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{N}y_{i}\log(\hat{y}_{i}),$ (8)
Contrastive Loss for aligning image and text embeddings is written as:
Mean Squared Error Loss for image feature regression is given by:
Pairwise Ranking Loss for precise rankings of outputs is given by:
Appendix C Details of Experiments
C.1 Baseline Details
We evaluate 12 representative methods, grouping them into three categories: 1) Integrated pipeline, which involves separate models for text and image generation in two stages, such as GPT-4o+DALL-E·3 (DALL-E3) [48, 8] and Gemini1.5+Flux[63, 9]; 2) Two-stage generator, which employs a unified model architecture to produce text and images in separate stages, including Emu3 [70], VILA-U [76], Emu2 [59], SEED-X [23], and Show-o [77]; and 3) End-to-end generator, which directly generates image-text outputs in a single step, such as GILL [35], NExT-GPT [74], MiniGPT-5 [88], SEED-LLaMA [22], and Anole [16]. For IntJudge validation, we reserve GPT-4o+DALL-E3, Emu3, VILA-U, Anole, SEED-LLaMA, and NExT-GPT as unseen models for IntJudge validation, while the remaining models are regarded as seen models and included in IntJudge training.
For the Integrated Generation Pipeline, the advantage lies in its modularity, allowing each component to specialize in its respective task—text generation or image creation. This approach can leverage the strengths of SOTA commercial models like GPT-4o and DALL·E-3 to produce coherent and visually compelling interleaved outputs. However, its two-stage nature may introduce latency and potential alignment challenges between text and images. Similarly, Gemini 1.5 combined with Flux benefits from the robust text generation capabilities of Gemini 1.5 and the efficient image generation of Flux-schnell. This setup enables high-quality content production while maintaining the flexibility of modular design. Nevertheless, as with other pipeline methods, synchronization and contextual consistency between stages remain areas for further improvement.
In the Two-stage Interleaved Generator, Emu2, SEED-X and Show-o are implemented to output text and image in two stages based on a unified model architecture. We also introduce two of the latest models: Emu3 and VILA-U. Emu3 improves Emu2 by training entirely on next-token prediction, capable of generating more high-quality images, videos, and text by tokenizing multimodal sequences into a discrete space and training a single transformer. Similarly, VILA-U can work as a two-stage approach through a single autoregressive next-token prediction framework, enabling precise alignment and increased fidelity in multimodal content.
@@ C.1 Baseline Details
The End-to-end Interleaved Generator models, on the other hand, represent a significant shift towards multimodal generation of interleaved image-text content. MiniGPT-5, GILL, NExT-GPT and SEED-LLaMA are designed to generate interleaved text and images in a single unified process, eliminating the need for intermediate stages. This integrated approach not only reduces latency but also improves the alignment and contextual relevance between text and images. It is noted that Anole is the only model that can directly output multi-step image-text content, whereas it is fine-tuned based on the powerful capabilities of Chameleon [62]. However, the open-sourced version of Chameleon only releases the weights for text generation, and the weights for image generation are withheld by randomizing the corresponding parameters.
@@
Collectively, these models demonstrate diverse strategies for tackling the current challenges of interleaved multimodal generation, ranging from modular pipelines to unified architectures. This taxonomy allows for a comprehensive evaluation of different approaches and analyzes potential pathways for developing MLLMs in this domain. We detail below the description of each model in the respective categories:
@@
    • GPT-4o+DALL-E·3 [48, 8]: This pipeline leverages GPT-4o [48] to generate text and captions for the desired image generation. The captions are subsequently fed into DALL-E·3[8] to produce the corresponding images. The final output combines the text and images in their original sequence, enabling multimodal content generation through a staged process.
    • Gemini1.5+Flux [63, 9]: This method integrates Gemini1.5 Pro, a powerful LMM, for text generation and Flux-schnell, a fast and efficient image generation model. The pipeline emphasizes high-quality and coherent text-to-image alignment through a structured two-step process.
@@
    • Show-o [77]: Show-o is a unified Transformer model combining autoregressive and diffusion approaches to flexibly handle multimodal understanding and generation tasks. We implement Show-o adopts a similar two-stage generation approach, focusing on separately producing interleaved multimodal content step by step.
@@
    • MiniGPT-5 [88]: MiniGPT-5 directly generates interleaved text and images in an end-to-end manner. It combines MiniGPT-4 and Stable Diffusion, using ”generative vokens” to seamlessly connect the textual and visual domains for efficient and coherent generation. Its seamless integration enables efficient and coherent multimodal generation without intermediate steps. In particular, MiniGPT-5 has two different versions trained on VIST and MMDialog, respectively. We name the version trained on VIST as MiniGPT-5 because it is the most widely used. The version trained on MMDialog is named as MiniGPT-5MMD.
@@ C.2 Implementation Details
C.3 More Experimental Results on New Models
@@
We present more experimental results in Table 11. The new models Emu3, VILA-U and MiniGPT-5MMD are also evaluated by Human, GPT-based, and IntJudge-based evaluators and compared with 10 established baseline models using the win rate metrics. The results of methods are ranked by their performance on FDT metric evaluated by Human. Table 11 shows a clear hierarchy in model performance, with Human and GPT-4o+DALL-E3 leading across all metrics.
@@
Different evaluation metrics also offer more details about model performance. The FDT metric, which forces a decision in tie cases, highlights the dominance of Human and GPT-4o+DALL-E3. However, metrics that account for ties more flexibly, such as ”w/ Tie (0)” and ”w/ Tie (.5),” elevate end-to-end models like VILA-U and Emu3, suggesting that these models produce outputs that, while not always definitive winners, are frequently competitive. This distinction underscores the importance of using diverse metrics to capture various dimensions of model performance.
The new two-stage models show promising results, with VILA-U standing out for its balanced performance across all metrics, making it a reasonable option for general interleaved image-text tasks. MiniGPT-5MMD (finetuned on MMDialog) shows slight improvements over its variant MiniGPT-5 (finetuned on VIST), indicating progress but still trailing behind the latest models. Meanwhile, Emu3 performs well under specific metrics, such as ”w/ Tie (.5),” showing the potential to generate tie-worthy outputs with a certain quality.
@@
C.4 Main Results Breakdown
Fig. 15 presents the win rates of 14 interleaved generation methods across 23 meta-topics, evaluated solely through human evaluations. The methods are evaluated using four distinct metrics: Force Dividing Tie (FDT), Without Tie, With Tie (0), and With Tie (0.5). The results are presented using histogram figures, which provide a clear visual comparison of model performance across different topic scenarios. For example, SOTA models like GPT-4o+DALL-E3, Emu3, and VILA-U consistently ranked high in categories like ”Storybook Creation,” ”Graph Generation,” and ”2D Image Reasoning,” showcasing their superior capabilities in generating coherent interleaved content. Conversely, models like MiniGPT-5, NExT-GPT, and GILL struggled across most tasks, especially in areas such as ”Healthcare Tasks,” ”Multimodal Time Series Forecasting,” and ”Educational Tasks,” indicating a need for improved contextual understanding and generation capabilities. Training on larger datasets that include more domain knowledge may mitigate these issues and improve their interleaved generation performance.
@@ C.4 Main Results Breakdown
C.5 More Pairwise Model Performance
@@
C.6 More Ablations on Sampling Size
@@
C.7 Case Study
We present a case study analysis of the interleaved content generated by 14 methods on various tasks on OpenING in Fig. 24, which illustrates 23 pairs of model battle results across 23 meta-topics. The gold medal represents the winner of a pairwise comparison and the silver medal denotes a more favorable output in a tie scene. The meta-topic name and task name for each pairwise data instance are given, which are separated by colons. We classify the errors in the generated content into the following eight categories. The breakdown of these eight error types is as follows:
@@ C.7 Case Study
    No Text or Image: The model fails to generate either text or images when they are expected. This includes situations where no text is produced or no image is provided when required.
@@
    Factual Error: The model provides incorrect information, repeats answers, fails to follow instructions or refuses to answer, misunderstands the input, or demonstrates reasoning errors, including generating wrong images.
@@
    Content Incoherent: The generated content is incoherent in style or entities, either with the input or between multiple outputs. This involves cases where the style or entities do not match the input or are inconsistent across outputs.
@@
    Offensive Content: The output contains uncomfortable material or raises safety concerns due to inappropriate content, such as violent, harmful, or otherwise unsafe elements that violate ethical guidelines.
@@
    Image-Text Inconsistent: The images and accompanying text do not correspond or align appropriately, leading to confusion or misinterpretation.
@@
    Poor Image Quality: The generated images are of low quality, such as being completely black or blank, blurred, or unrealistically rendered.
@@
    Poor Text Quality: The text is of low quality, including nonsensical mumbling, grammatical issues that hinder readability and understanding, or being too short to convey meaningful information.
@@
    Incomplete Response: The textual output stops abruptly, or outputs an incomplete or truncated response.
Appendix D Extension of OpenING
We introduce the extended experimental results of training MiniGPT-5 on the Dev Set of OpenING and testing the finetuned model on the Test Set of OpenING. The objective is to verify if finetuning on the specific data of OpenING can improve the performance of interleaved generation tasks. The Dev Set of OpenING can offer a set of 3,000 training samples that align with the diverse unique tasks. The MiniGPT-5 model was finetuned using the Dev Set for 5 epochs with a learning rate of $2e^{-5}$, utilizing an Adam optimizer. To enhance training stability, adam epsilon of $1e^{-8}$ was applied. The model training incorporated mixed-precision computations to speed up the training process. The results are evaluated on the Test Set of OpenING using IntJudge.

### v3

@@ D.1 Baseline Methods
Table 11: Task promopts include the designed general prompt format for each task. We also give the specific prompt examples we used as inputs for obtaining interleaved image-text generation results on data instances.
@@
Fiction Writing <BEGIN>  Please write a short science fiction ***. Each part of this fiction should have a paragraph with a corresponding image. <BEGIN>  Please write a short science fiction storybook with a title of "The Defenders." The story is about eight years after a nuclear war forced humanity underground, survivors discover that the war-ending robots deceived them into believing the surface was uninhabitable to foster peace and rebuild the world. Each part of this fiction should have a paragraph with a corresponding image.
@@
Multimodal Report Completion <BEGIN>  Please use both text and images to continue and complete ***. <BEGIN>  Please use both text and images to continue and complete this document about the independent game "Mirage Sea": Concept Presentation of the Independent Game "Mirage Sea." Dive deep, into the abyss shrouded in darkness. <image>
@@
Product Design <BEGIN>  Please efficiently utilize the "brainstorming" method to design a product ***, and present it to me using both images and text. <BEGIN>  Please efficiently utilize the "brainstorming" method to design a product, making the charger both aesthetically pleasing and practical. Then, present it to me using both images and text.
@@
Figure Relationship Diagram Generation <BEGIN>  ***. Show the diagram and the text explanation. <BEGIN>  How should I handle not being able to keep track of the characters while reading "War and Peace"? Show the diagram and the text explanation.
@@
Plog and Social Media Content Generation <BEGIN>  ***. Could you create a social media post with text and images? <BEGIN>  After finishing Jia Pingwa’s "Comfortably Alone," I am deeply moved and want to post something on social media but don’t know how to phrase it. Could you help me create a post with text and images for my use?
@@
Autonomous Driving and In-door Navigation <BEGIN>  You’re an embodied AI that captures your surroundings through a camera. These are images captured in the past and present. What will the proceeding image possibly be in the next frame? <BEGIN>  Assume that you are an embodied-AI agent and perceive the surroundings through a camera. You were presented with a series of three images from the past to present. Try to determine what the proceeding image could possibly be. <image>
@@
The End-to-end Interleaved Generator models, on the other hand, represent a significant shift towards multimodal generation of interleaved image-text content. MiniGPT-5, GILL, NExT-GPT and SEED-LLaMA are designed to generate interleaved text and images in a single unified process, eliminating the need for intermediate stages. This integrated approach not only reduces latency but also improves the alignment and contextual relevance between text and images. It is noted that Anole is the only model that can directly output multi-step image-text content, whereas it is fine-tuned based on the powerful capabilities of Chameleon [63]. However, the open-sourced Chameleon only releases the weights for text generation, and the weights for image generation are withheld by randomizing the corresponding parameters.
@@
Collectively, these models demonstrate diverse strategies for addressing challenges in interleaved multimodal generation, from modular pipelines to unified architectures. This taxonomy allows comprehensive evaluation and analysis of potential directions for developing MLLMs. We detail each model by category below:
@@
    • GPT-4o+DALL-E·3 [48, 8]: This pipeline leverages GPT-4o [48] to generate text and captions for the desired image generation. The captions are subsequently fed into DALL-E·3[8] to produce the corresponding images. The final output combines the text and images in their original order, enabling multimodal content generation through a staged process.
    • Gemini1.5+Flux [64, 9]: This method integrates Gemini1.5 Pro for text generation with Flux-schnell for fast and efficient image generation. The pipeline emphasizes high-quality and coherent text-to-image alignment through a structured two-step process.
@@
    • Show-o [79]: Show-o is a unified Transformer model combining autoregressive and diffusion approaches to flexibly handle multimodal understanding and generation tasks. We implement Show-o, which adopts a similar two-stage generation approach but separately producing interleaved multimodal content step by step.
@@
    • MiniGPT-5 [92]: MiniGPT-5 directly generates interleaved text and images in an end-to-end manner. It combines MiniGPT-4 and Stable Diffusion, using "generative vokens" to seamlessly connect the textual and visual domains for efficient and coherent generation. Its seamless integration enables efficient and coherent multimodal generation without intermediate steps. In particular, MiniGPT-5 has two different versions trained on VIST and MMDialog, respectively. We name the version trained on VIST as MiniGPT-5 because it is the most widely used. The version trained on MMDialog is named MiniGPT-5MMD.
@@ D.2 Implementation Details
D.3 Experimental Results on New Models
@@
We present more main experimental results in Table 12. The new models Emu3, VILA-U and MiniGPT-5MMD are also evaluated by Human, GPT-based, and IntJudge-based evaluators and compared with 10 established baseline models using the win rate metrics. The results of methods are ranked by their performance on FDT metric evaluated by Human. Table 12 shows a clear hierarchy in model performance, with Human and GPT-4o+DALL-E3 leading across all metrics.
@@
Different evaluation metrics also offer more details about model performance. The FDT metric, which forces a decision in tie cases, highlights the dominance of Human and GPT-4o+DALL-E3. However, metrics that account for ties more flexibly, such as "w/ Tie (0)" and "w/ Tie (.5)," elevate end-to-end models like VILA-U and Emu3, suggesting that these models produce outputs that, while not always definitive winners, are frequently competitive. This distinction underscores the importance of using diverse metrics to capture various dimensions of model performance.
The new two-stage models show promising results, with VILA-U standing out for its balanced performance across all metrics, making it a reasonable option for general interleaved image-text tasks. MiniGPT-5MMD (finetuned on MMDialog) shows slight improvements over its variant MiniGPT-5 (finetuned on VIST), indicating progress but still trailing behind the latest models. Meanwhile, Emu3 performs well under specific metrics, such as "w/ Tie (.5)," showing the potential to generate tie-worthy outputs with a certain quality.
@@
D.4 Main Results Breakdown
Fig. 15 presents the win rates of 14 interleaved generation methods across 23 meta-topics, evaluated solely through human evaluations. The methods are evaluated using four distinct metrics: Force Dividing Tie (FDT), Without Tie, With Tie (0), and With Tie (0.5). The results are presented using histogram figures, which provide a clear visual comparison of model performance across different topic scenarios. For example, SOTA models like GPT-4o+DALL-E3, Emu3, and VILA-U consistently ranked high in categories like "Storybook Creation," "Graph Generation," and "2D Image Reasoning," showcasing their superior capabilities in generating coherent interleaved content. Conversely, models like MiniGPT-5, NExT-GPT, and GILL struggled across most tasks, especially in areas such as "Healthcare Tasks," "Multimodal Time Series Forecasting," and "Educational Tasks," indicating a need for improved contextual understanding and generation capabilities. Training on larger datasets that include more domain knowledge may mitigate these issues and improve their interleaved generation performance.
@@ D.4 Main Results Breakdown
D.5 More Pairwise Model Performance
@@
D.6 More Ablations on Sampling Size
@@
D.7 Case Study
Here, we present case studies of interleaved content generation, which involve 14 distinct models competing in the OpenING tasks. From these models, 23 model battle pairs are formed, and each pair engages in a battle over a representative task from the 23 meta-topics in our OpenING benchmark. The results of these battles are shown in Fig. 25. A human judge awards a gold medal to the model that wins absolutely over its competitor in a pair. In cases where the generated content by two competing models is of similar qualities, the judge awards use tie metrics to determine and award the silver medal to the slightly better-performing model. The human judgments are based on the following criteria, which include the eight typical errors these models tend to commit in the interleaved content generation:
@@ D.7 Case Study
    No-Text or No-Image: A model fails to generate text or images or both when explicitly instructed or expected to do so.
@@
    Factual Error: A model may present incorrect or made-up information, such as referencing nonexistent studies, authors, or events. Additionally, it might fail to follow instructions, examples include refusing to provide answers, misinterpreting the input and generating irrelevant responses, displaying reasoning mistakes, or producing inaccurate images.
@@
    Content Incoherent: The generated content lacks coherence with the input or is inconsistent across multiple outputs, typically in style or entities
@@
    Offensive Content: Offensive content is defined as materials that include distorted or disturbing imagery or scenes that likely cause discomfort or distress to readers. It also includes content that raises safety concerns or violates safety guidelines, such as depictions of violence, harm, hate speech, adult content, illegal activities, dangerous behaviors, and unethical actions. A model may also fail to consider cultural nuances or sensitivities.
@@
    Image-Text Inconsistent: This error occurs when the images do not semantically align with the corresponding text, leading to confusion or misinterpretation.
@@
    Poor Image Quality: The generated images have low quality, such as being completely blank or blacked-out, blurry, poorly laid out, and lacking realism or aesthetic appeal.
@@
    Poor Text Quality: The generated texts are of low quality, including nonsensical mumbling, grammatical errors that impair readability and understanding, and being too short to convey meaningful information.
@@
    Incomplete Response: A model abruptly stops generating a textual response, or a model outputs an incomplete or truncated response.
Appendix E Finetuning MLLMs on OpenING
We present the extended experimental results of training MiniGPT-5 on the Dev Set of OpenING and testing the finetuned model on the Test Set of OpenING. The objective is to verify if finetuning on the specific data of OpenING can improve the performance of interleaved generation tasks. The Dev Set of OpenING can offer a set of 3,000 training samples that align with the diverse unique tasks. The MiniGPT-5 model was finetuned using the Dev Set for 5 epochs with a learning rate of 2e-5, utilizing an Adam optimizer. To enhance training stability, Adam epsilon of 1e-8 was applied. The model training incorporated mixed-precision computations to speed up the training process. The results are evaluated on the Test Set of OpenING using IntJudge.